Circular No. 6570

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 960720 J. Greiner, Astrophysical Institute, Potsdam; and J. Heise, Space Re-

search Organization of the Netherlands, communicate: “The brightest of the nine x-ray sources detected with the ROSAT HRI on 1996 Aug. 31 (IAUC 6487), RX J1730.7+4906, is located within the revised error circle of GRB 960720 (IAUC 6569). It coincides with QSO 1729+491 (= 4C 49.29), which is a strong radio source at z = 1.038 (Walsh et al. 1984, MNRAS 211, 105). Its HRI countrate of 0.0041 ± 0.0009 count/s corresponds to an unabsorbed x-ray luminosity of 5 × 10^{44} erg/s in the range 0.1–2.4 keV (assuming a power-law spectrum of photon index 1.9, corrected for galactic absorption). The likelihood for the positional coincidence of a radio-loud AGN of that x-ray and radio brightness with the 3′ GRB error box is ∼ 2 × 10^{−4} (Brinkmann et al. 1995, A.Ap. Suppl. 109, 147).”

L. Piro, E. Costa, M. Feroci, and M. Cinti, Istituto di Astrofisica Spaziale, Consiglio Nazionale delle Ricerche (CNR), Frascati; F. Frontera, D. Dal Fiume, L. Nicastro, M. Orlandini, and G. Pizzichini, Istituto Tecnologie e Studio Radiazioni Extraterrestri, CNR, Bologna; J. Heise and R. Jager, Space Research Organization of the Netherlands, Utrecht; F. Fiore, P. Giommi, A. Antonelli, and J. M. Muller, BeppoSAX Science Data Center, Rome; A. Parmar, Space Science Department, ESTEC, The Netherlands; M. C. Maccarone, Istituto Fisica Cosmica e Applicazioni Informatica, CNR, Palermo; L. Chiappetti, Istituto di Fisica Cosmica e Tecnologie Relative, CNR, Milan; and R. C. Butler, Agenzia Spaziale Italiana, Rome, report: “We analyzed data from the BeppoSAX follow-up observation of the Wide-Field-Camera field of GRB 960720 (IAUC 6467, 6480) made on Sept. 3.715–4.708 UT. Within the updated error region of this GRB (IAUC 6569), we detect one source with a probability of chance occurrence of ∼ 1 × 10^{−4}. The estimated x-ray flux (2–10 keV) is (1.0 ± 0.3) × 10^{−13} erg cm^{−2} s^{−1}, and the source is located at α = 17^{h}30^{m}45^{s}, δ = +49^{o}06′.1 (equinox J2000.0; conservative error radius of 50′′); this position is coincident with that of a bright radio-loud QSO, 4C 49.29, which we identify as the counterpart of the x-ray source. The probability that such a (radio) bright QSO is found by chance within the WFC error box is < 2 × 10^{−4}. However, we note that this is the only case so far reported of an object of this type being associated to a small (a few arcmin^{2}) GRB error box.”

1997 March 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green